### v1

# Multi-perspective Feedback-attention Coupling Model for Continuous-time Dynamic Graphs

Abstract
Recently, representation learning over graph networks has gained popularity, with various models showing promising results. Despite this, several challenges persist: 1) most methods are designed for static or discrete-time dynamic graphs; 2) existing continuous-time dynamic graph algorithms focus on a single evolving perspective; and 3) many continuous-time dynamic graph approaches necessitate numerous temporal neighbors to capture long-term dependencies. In response, this paper introduces the Multi-Perspective Feedback-Attention Coupling (MPFA) model. MPFA incorporates information from both evolving and raw perspectives, efficiently learning the interleaved dynamics of observed processes. The evolving perspective employs temporal self-attention to distinguish continuously evolving temporal neighbors for information aggregation. Through dynamic updates, this perspective can capture long-term dependencies using a small number of temporal neighbors. Meanwhile, the raw perspective utilizes a feedback attention module with growth characteristic coefficients to aggregate raw neighborhood information. Experimental results on a self-organizing dataset and seven public datasets validate the efficacy and competitiveness of our proposed model.
Index Terms:
Continuous-time dynamic graphs, Multi-perspective, Feedback attention, Evolving and raw perspectives, Representation learning
publicationid: pubid: $0$–$00$ 2021 IEEE
Figure 1: A toy example illustrates a continuous-time dynamic graph with multiple complex dynamics, where $t_{0}$-$t_{3}$ represent timestamps. The evolution of the graph from $t_{0}$ to $t_{3}$ captures the entire process, revealing that new interactions between nodes are established at each timestamp. At $t_{1}$, a new node joins the network and interacts with another node (red edge). At $t_{2}$, co-occurrence events occur (green edges), and at the final timestamp, the connection between nodes 1 and 6 is re-established (blue edge). Conversely, a discrete-time dynamic graph can be formed by extracting snapshots (e.g., snapshots $t_{0}$ and $t_{2}$) from the temporal network. If only the final state of the graph evolution is considered, a static graph can be obtained.
I Introduction
Representation learning has illuminated the analysis of graph-structured data, leveraging its powerful capability to encode structures and properties of graph networks. Typically, it acquires node embedding representations by transforming high-dimensional and non-Euclidean graphs into lower-dimensional vector embeddings [1]. These node embeddings are then employed in various graph analysis tasks, including graph classification, edge prediction, or node classification [2, 3]. While existing representation learning methods have demonstrated promising performance across numerous domains [4, 5], several challenges remian. The first issue with the majority of graph representation learning methods is that they primarily focus on static graph networks or discrete-time dynamic graphs[6, 7, 8]. In reality, graph data in most domains (e.g., natural language processing, social networks, recommendation systems) exhibit complex dynamics, meaning that the network structures and properties constantly change over time (In Figure 1, we illustrate a straightforward process depicting changes in a dynamic network). The second problem is that existing continuous-time dynamic graph algorithms only consider the evolving perspective, ignoring other aspects of the graph. Finally, most continuous-time dynamic graph approaches require a large number of temporal neighbors to capture long-term dependencies, which can be computationally expensive. Models for continuous-time dynamic graph representation learning need to support the addition of nodes and edges, as well as feature changes, and generate node embeddings at any time. While the usage of static graph methods (e.g., GAT [8] and GraphSAGE [9]) on temporal networks is possible in some cases, taking no account of the temporal information can severely weaken the ability to capture the crucial insights of networks. Therefore, there is a need for the development of more and more effective models for these dynamically changing data.
Existing techniques for learning representations on dynamic graphs can be divided into two folds: discrete-time dynamic graphs and continuous-time dynamic graphs. Discrete-time dynamic graphs are represented as a sequence of network snapshots at specific time intervals [10, 11, 12]. Since the network is represented in snapshots, the static graph analysis method can be used for each snapshot with a fixed topology. The repeated static method and the time information learning between the snapshot embeddings can collectively give insight into the graph’s dynamics. The existing methods of discrete dynamic graphs typically include stacked structures, integrated structures, graph auto-encoders, and generative models [13, 14, 15, 16, 17, 18]. However, such dynamic graph representation makes a model sensitive to the snapshot sampling interval. When the sampling interval is too large, much information in the network will be lost, leading to the model only capturing limited evolutionary information. Recently, several methods have been proposed for supporting the continuous-time settings[19, 20, 21]. These methods view dynamic graphs as a stream of events occurring in chronological order, so their typical processing methods follow an event-level paradigm. Examples include temporal random walk-based methods[22, 23], RNN-based methods[24, 25], attention-based methods[26], and transformer-based methods[27, 28]. We find that these methods need to look up a large number of historical interactions in order to obtain long-term dependencies[29, 30, 28]. This leads to increased computational and time costs. When the acquired historical interactions are short, the performance of the model degrades. In addition, these methods focus only on a single evolving perspective. Although memory-based methods can use shorter historical interactions to acquire long-term dependencies, they are also limited by the singularity of the perspective, which prevents the model from learning the nature of the raw interactions and the growth characteristics from the raw interactions to the current state.
In this paper, we propose corresponding solutions for each of the above challenges: For the first challenge, our proposed MPFA can effectively complement the continuous-time dynamic graph model. For the second challenge, due to the time-varying nature of continuous-time dynamic graphs, when a model aggregates the temporal neighbor information of a node, these neighbors may no longer be in the interaction initial state (they may have interacted with other nodes multiple times). In many application scenarios, only considering evolutionary interactions is often insufficient and may lead to information redundancy or inference errors. (For example, in the shopping system, after a person has purchased a mobile phone, his purchase of a mobile phone protective case is generally related to the purchase of a mobile phone and should not be affected too much by other events). Therefore, we propose to preserve the raw state to improve the performance and generalization of a model. First, we design an evolving perspective that can capture the real-time dynamics of the nodes at any time. All participating nodes in this perspective contain the latest evolutionary content of the network. To distinguish the importance of different neighbors, we use temporal self-attention to discriminatively aggregate the neighbor information. Second, we design a raw perspective. The raw interaction represents the past facts and reveals the essence of events. When the network returns to observe these events after learning, it will have deeper insights. We achieve the information fusion of the raw interactions by using the feedback attention coefficients with model growth characteristics. Finally, the two perspectives learn interactively, instructing each other to obtain the final target node embedding. For the third challenge, to achieve the property of capturing long-term dependencies using a small number of temporal neighbors, we perform a dynamic update of the evolving information in the evolving perspective immediately after each event. This allows a form of compressed coding to record the historical events for each node. Overall, the main contributions of this paper are as follows:
  • We propose a novel inductive model called MPFA, working on continuous-time dynamic graphs described as an ordered sequence of timepoint events.
  • We make the first attempt to explore continuous-time dynamic graphs from raw and evolving perspectives and design different attentions for them. The two views enable our model perform better in inductive prediction tasks as well as in dynamic classification tasks.
  • We design a feedback attention coefficient with model growth characteristics to learn the neighborhood information closely related to the target node with deeper insights. In addition, dynamically updating evolving information can be used to capture long-term dependencies using a small number of temporal neighbors.
  • Experimental results on 1 self-organized dataset and 7 public datasets show that our proposed model can achieve better results in both dynamic link prediction and dynamic node classification tasks.
II Related work
In this section, we briefly introduce the related work on static and discrete-time dynamic graphs and then elaborate on the existing work on continuous-time dynamic graphs.
Static graphs: The purpose of static graph representation learning is to perform node embedding, or sub-graph embedding [1, 2, 3]. The approaches based on the spectral domain apply the convolution kernel on the input signal in the spectral space and then use the convolution theorem to aggregate the information between nodes [31, 32, 33]. The core of graph convolution of the spectral domain is to operate the Laplacian Matrix by projecting, expanding, or approximating. Since this kind of method requires a fixed topological structure, it cannot be applied to the changing temporal graph. The definition of the spatial domain methods is more flexible[34] and can be extended to large-scale graphs. It mainly selects the relevant nodes as the neighborhood and then aggregates the neighborhood information under the constraint of the adjacency matrix (e.g., GraphSAGE [9], GAT [8], and PGC [35]). These methods have achieved good performance when processing static graphs. However, the performance will be poor when operating on temporal graphs due to the lack of consideration of temporal information.
Discrete-time dynamic graphs: Early modeling approaches for dynamic graphs mainly focused on discrete-time methods[11, 12, 36]. Since the dynamic graph is represented as an ordered sequence of snapshots, one of the advantages of this formation is that static graph models can be utilized on each snapshot. The most straightforward way for such discrete approaches is to use a static Graph Neural Network (GNN) to obtain the embedding of each snapshot. Then the sequence (e.g., RNN or Lstm) or attention model is utilized for encoding the temporal dependence between snapshots [11, 14, 15]. Simultaneous modeling of topological structure and temporal information is another solution, and it can also be considered as an encoder [12, 16]. It mainly combines the static graph method with the sequence model to form a layer so that the temporal dependence and network structure can be captured simultaneously in one layer. In addition, there are some other approaches, including matrix factorization, autoencoders, and generative models[17, 37, 36, 18].
Figure 2: The Interaction between raw and evolving information (i.e., transmission of information on a timeline). The model learns node embedding based on the two informations and predicts whether event (1, 6) occurs at time $t$. For node 1 (similar to node 6), the dotted arrow represents the influence of evolving information on the raw information, and vice versa. The blue and red arrows mean the effect of the raw and evolving neighborhood on the current event; The effect of the node’s growth information on the raw ation is indicated by the orange dotted arrows. The purple dotted arrows represent that the evolving neighborhood has an impact on its own raw information.
Continuous-time dynamic graphs: Processing continuous-time dynamic graphs is a challenging task, and only recently, some approaches have been proposed [38]. CTDNE[22] modifies the path sampling method of DeepWalk to make the path meet the actual occurrence sequence according to time, thereby making the sampling path more reasonable. Based on LSTM, the architecture of DyGNN[19] includes two components: An update module updating the embedding of the nodes involved in an event; A propagation module responsible for propagating the updated state to the node neighbors. JODIE [24] employs two RNNs to design a coupled recurrent neural network for node embedding, containing three modules: update, projection, and future interaction prediction. Some methods are based on variants of random walks[20, 23], which integrate continuous-time information by constraining on transition probability matrix. TGAT[26] takes the self-attention mechanism as a building block and develops a novel functional temporal encoding technique according to the classical Bochner theorem of harmonic analysis. DYREP[21] considers two dynamic processes on the temporal network: the Association Process and the Communication Process. The former represents changes in network topology, and the latter denotes changes in dynamics on the network. TGN [25] designed a node memory module to store the long-term dependencies of nodes. CAWN [29] initially generates a multitude of causal anonymous walks for each node. Subsequently, it encodes each walk using a recurrent neural network and combines these walks to attain the ultimate node representation. EdgeBank [30] is a purely memory-based method with no trainable parameters for transductive dynamic prediction tasks. TCL [27] uses a breadth-first search algorithm to generate interaction sequences of nodes and a graph transformer to obtain node embedding representations. DyGFormer[28] uses Transformer as the encoder. Neighbor co-occurrence coding as well as patching techniques are used in the sequence processing scheme. GraphMixer [39] uses a time encoding that requires no training and uses MLP-Mixer as the encoder to obtain node embedding representations.
Figure 3: The overall architecture of MPFA. $t_{p_{1}}-t_{p_{6}}$ denote the interaction time before time $t$. The evolving perspective (bottom) aggregates information from constantly evolving neighborhood by using a temporal self-attention module; For the raw perspective (top), the feedback attentions with growth characteristics calculated by the Feedback Coefficient module are utilized to aggregate the raw neighborhood information; In the above two processes, the Raw and Evolving Information Preserved modules play a critical role. Then the acquired evolving and raw features are coupled to teach each other through the Attention Coupling module for final temporal graph embedding. After completing the learning of the two perspectives (gray arrow process), MPFA performs the update of EIP-RIP (red arrow process).
III Definations
III-A Continuous-time dynamic graphs
A continuous-time dynamic graph is described as a sequence of time point events, where each event involves two nodes at a certain time. Formally, it can be represented as $G(T)=(V(T),E(T))=[x({t_{0}}),\;x({t_{1}}),\;x({t_{2}}),\>...]$. Given any time $t$, it can be denoted as $G(t)=(V(t),E(t))=[x({t_{0}}),x({t_{1}}),...,\>x({t})]$, where $V(t)$ and $E(t)$ are sets of nodes and edges, respectively. $x(t)=(v_{i},v_{j},e_{i,j}(t))$ is an event happened at time $t$, where $v_{i}$ and $v_{j}$ mean the nodes, $i$ and $j$ are the node numbers. $e_{i,j}(t)$ denotes the edge feature. It should be noted that a pair of nodes $(v_{i},v_{j})$ may build multiple interactions at different time points.
III-B Continuous-time dynamic graph representation learning
The goal of Continuous-time dynamic graph representation learning is to learn node representations and use them for downstream tasks (e.g., dynamic node classification, and future link prediction). Given a temporal graph $G$, temporal graph embedding aims to seek a mapping function $f:V\longrightarrow\mathbb{R}^{d}$, where $d$ is the dimension of node representation. At the time $t$, the temporal graph embedding can be denoted by $Z(t)=({z_{1}}(t),...,{z_{n}}(t))$.
III-C Raw and evolving information preserved
According to the evolution direction of the network in time, we can divide the temporal events of a node into two states, the raw and evolving state. As shown in Figure 2, the raw state maintains the essence of an event at that time, and the evolving state is a description of a node’s ongoing change.
Evolving information preserved: Different from conventional approaches, we design a non-parameterized vector for each node to preserve the evolving information instead of directly parameterizing the node embedding. The evolving information of node $v_{i}$ is denoted as ${h_{i}}(t)$ at time $t$, representing all events of the node that occurred before time t in a compressed encoding form. This encoding captures long-term dependencies while utilizing only a small number of temporal neighbors. During the network evolving, it is updated after each event. When an event $x(t)$ happens between node $v_{i}$ and $v_{j}$, we first obatin the current event information ${X_{i,j}(t)}=W_{evol}[{z_{i}}({t^{-}})||{e_{i,j}}(t^{-})||{z_{j}}({t^{-}})||\Phi(\Delta t)]$, and then update the evolving information of $v_{i}$ as follows:
${h_{i}}(t)=update({X_{i,j}(t)},\;{h_{i}}({t^{-}})),$ (1)
where $\Phi(\cdot)$ denotes a generic time embedding [40, 26], and ||||| | represents a connection symbol. ${t^{-}}$ is the last interaction time of node $v_{i}$. $W_{evol}$ is the parameter used to scale the dimension. $z_{i}(t^{-})$ and $z_{j}(t^{-})$ denote the last embedding representation of nodes $v_{i}$ and $v_{j}$, respectively. Update(.) means a learnable function, which can be GRU or RNN. When a new node appears in the network, its evolving information will be initialized to a vector with zero values.
Raw information preserved: Due to the time-varying nature of continuous-time dynamic networks, each node will change (e.g., interacting with other nodes) over time. Thus, the nodes would no longer maintain the original interactive state. It is essential to keep the initial event state between node $v_{i}$ and $v_{j}$ unchanged so that the model can better understand the nature of historical events. It can be considered a growth process from the raw state to the evolving state. Given the raw events $[(v_{i},v_{1},e_{i,1}(t_{1})),...,(v_{i},v_{j},e_{i,j}(t_{j}))]$ of $v_{i}$, their raw event information is represented as $[{r_{1}}({t_{1}}),...,{r_{j}}({t_{j}})]$. How to design the raw information of a node is worth exploring, which will also be one of the directions of our future work. Here, for convenience, We denote the raw information as ${r_{j}(t_{j})}=W_{raw}[{z_{j}}({t^{-}_{j}})||{e_{i,j}}(t_{j}^{-})||{z_{i}}({t^{-}_{j}})||\Phi(\Delta t_{j})]$.
IV Representation learning on raw and evolving perspectives
Given the problem definitions above, we describe our proposed model in this section. We first illustrate the general framework of the MPFA in Figure 3 and then elaborate on its composition. MPFA, a multi-perspective feedback-attention coupling model, can project the temporal graph $G(t)$ at time t into low-dimensional space to generate node embeddings. The evolving perspective of MPFA consists of a temporal self-attention module aggregating constantly evolving temporal neighborhood information. In this evolving perspective, dynamically updated evolutionary information can capture long-term dependencies using fewer temporal neighbors. The raw perspective is an feedback attention for generating the aggregation of the raw neighborhood. To implement the feedback attention, the feedback attention coefficients with model growth characteristics are computed by a feedback module. Then, the two perspectives learn collaboratively to enhance the generalization and prediction capabilities of the model. More specific details of each module are discussed in the following subsections.
IV-A Evolving perspective
In the evolving perspective, a sefl-supervised temporal attention is designed to obtain the evolving neighborhood embedding representation of node $v_{i}$ at time t. This perspective is used to capture the real-time dynamics of the network. Given a node $v_{i}$ involved in an current event, its evoleving information at time $t$ and just before $t$ are denoted by ${h_{i}}({t})$ and ${h_{i}}({t^{-}})$, respectively. We denote by ${N_{i}}(t)=[{v_{1}},...,{v_{j}}]$ the temporal neighbors, and $[{h_{1}}(t^{-}),...,{h_{j}}(t^{-})]$ the evolving representation of neighbors just before time $t$. For the temporal self-attention, the inputs are ${h_{i}}({t^{-}})$, the current timepoint t, $[{h_{1}}(t^{-}),...,{h_{j}}(t^{-})]$, and the edge features $[{e_{i,1}}({t^{-}}),...,{e_{i,j}}({t^{-}})]$. The specific calculation process is as follows:
$\begin{split}head&=atten(Q(t),K(t),V(t))\\ &=Soft\max(\frac{{Q(t)K{{(t)}^{T}}}}{{\sqrt{d}}})V(t)\end{split}$ (2)
$P_{N_{i}(t)}^{e}=Concat(hea{d_{1}},hea{d_{2}}){W_{con}}$ (3)
The query $Q(t)={h_{i}}({t^{-}})||\Phi(0)$ is the relevant information of the target node. The key $K(t)=[{h_{1}}(t^{-})||{e_{i,1}}({t^{-}})||\Phi(t-t_{1}),\;.\;.\;.\;,{h_{j}}(t)||{e_{i,j}}({t^{-}})||\Phi(t-{t_{j}})]$ and value $V(t)=K(t)$ are the relevant information of temporal neighbors of the target node.
IV-B Raw perspective
IV-B1 Feedback coefficients with growth characteristics
For the raw perspective, we consider the factor that it is a growth process from the raw to the evolving state. After the model has grown over a period of time, the model will be more profound insights into the previous events. This insight is reflected not only in the feedback coefficients but also in the evolution information of the neighbors, which the model uses to obtain more useful information for future events from the raw interaction. In addition, During the learning process of node representation, historical interactions that contribute to the current event should not be too much disturbed by other unrelated matters. First, we compute feedback coefficients with growth properties, and then use the feedback coefficients to aggregate the raw interaction information. Given the embedding information $z_{i}(t^{-})$ and $z_{j}(t^{-})$ of nodes $v_{i}$ and $v_{j}$ just before time $t$, we compute the feedback coefficients from the time $t_{j}$ when $v_{i}$ and $v_{j}$ interacted to the present time $t$.
${I_{i,j}}(t_{j}\to t)=\psi({g_{i,j}}({t_{j}\to t})),$ (4)
where ${g_{i,j}}({t_{j}\to t})$ is used to learn the growth properties from time $t_{j}$ to $t$. Here we use only simple linear mappings, which can be computed by:
${g_{i,j}}(t_{j}\to t)=act({W_{r_{0}}}[{h_{i}}(t^{-})||{h_{j}}(t^{-})||\Phi(t-{t_{j}})]).$ (5)
Here, ||||| | is the concatenation operator, and act(.) is the active function. We stipulate that the feedback coefficient must be positive. Therefore, the external function must satisfy that its values are all greater than or equal to zero. For simplicity, we so we choose sigmoid as the external function.
${I_{i,j}}(t_{j}\to t)={(1+{e^{W_{r_{1}}{g_{i,j}}({t_{j}\to t})}})^{-1}}$ (6)
The final feedback coefficients can be obtained as follows:
${a_{i,j}}(t_{j}\to t)=\frac{{\exp({I_{i,j}}({t_{j}\to t}))}}{{\sum\nolimits_{j^{{}^{\prime}}\in{{\rm N}_{i}}(t),}{\exp({I_{i,j^{{}^{\prime}}}}({t_{j}\to t}))}}}.$ (7)
These coefficients are then utilized to aggregate the raw neighborhood information for target node i.
$P_{N_{i}(t)}^{r}=\sum\limits_{j\in{N_{i}}(t)}{{a_{i,j}}(t_{j}\to t)*{W_{trans}}\;}{C_{j}}({t_{j}}),$ (8)
where ${C_{j}}({t_{j}})=[{r_{j}}({t_{j}})||\Phi(t-{t_{j}})]$, and $W_{trans}$ is the linear transformation parameter; two linear transformations are used here. We have now finished presenting information capture from both perspectives, and the following section describes the interrelated learning of these two perspectives.
IV-C Attention coupling learning
As raw and evolving perspectives mutually drive the evolution of the temporal network, which alternately influence the learning process of network embeddings. In order to enable the model to have better performance, the evolving and raw embeddings are coupled to guide each other, which can also be regared as the interaction between the two attentions. First, we consider three types of information, i.e., changes in the node itself, raw information, and evolving information:
$z_{i}^{1}(t)={W_{node}}[{h_{i}}({t^{-}})||\Phi(t-t^{-})],$ (9)
$z_{i}^{2}(t)={W_{r}}[{h_{i}}({t^{-}})||P_{N_{i}(t)}^{e}],$ (10)
$z_{i}^{3}(t)={W_{e}}[{h_{i}}({t^{-}})||P_{N_{i}(t)}^{r}],$ (11)
Then we pass it to a two-layer feedforward neural network (FNN) for attention coupling learning.
${z_{i}}(t)=FNN([z_{i}^{1}(t)||z_{i}^{2}(t)||z_{i}^{3}(t)])$ (12)
After obtaining the embedding of nodes $v_{i}$ and $v_{j}$ at time t, the MLP decoder can be used to determine whether nodes $v_{i}$ and $v_{j}$ are interacting at time $t$, i.e., $p_{i,j}(t)=MLP(z_{i}(t),z_{j}(t))$.
IV-D Training
The optimization of parameters typically depends on downstream tasks. Here link prediction task is chosen to train MPFA. Given N events arranged in chronological order, we randomly sample the same number of negative pairs from dataset that not occur in the batch. For the sake of simplicity, after a batch of events is decoded , the Binary Cross Entropy function ${l_{n}}=-({y_{n}}*\log(\delta({s_{n}})))+(1-{y_{n}})*\log(1-\delta({s_{n}})),n\in 2N$ is leveraged to train the parameters, where $y_{n}$ denotes the label of the $n^{th}$ sample pair, and the $s_{n}$ represents the probability that the $n^{th}$ sample pair is predicted to be positive. The $\delta\left({{s_{n}}}\right)$ is the sigmoid function for numerical stability.
V Experiments
To validate the effectiveness of the model, we experiment extensively on a self-organizing dataset and seven public datasets. We compare it to baselines in dynamic link prediction and dynamic node classification, using both transductive and inductive inference. In addition, we perform model ablation, hyperparameter analysis, attention coefficient visualization, time consumption analysis, and long-term dependence analysis.
V-A Datasets
UNVote: [^1] https://zenodo.org/records/7008205#.Yv_a_3bMJPZ UNVote is the roll-call voting network of the United Nations General Assembly. It consists of 201 nodes and 1,035,742 edges.
LastFM: [^2] https://snap.stanford.edu/jodie/ This dataset is a bipartite network consisting of interactions between users and songs. The network contains about 1,000 users and 1,000 of the most listened to songs, and these nodes have generated a total of 1,293,103 interactions.
SocialEvo:11footnotemark: 1 This is a university community cell phone proximity monitoring network that tracks student activity over an eight-month period, totaling 2,099,519 edges.
MOOC:22footnotemark: 2 MOOC is a network of online course interactions whose nodes are students and course content. Each interaction represents a student’s access to the course. The dataset has a total of 7144 nodes and 411,749 edges.
Reddit:22footnotemark: 2 In this dataset, the nodes are 10,000 active users and 1,000 active subreddits, and an interaction is a post written by users on subreddits, resulting in 672,447 interactions. Dynamic labels represent whether a user is banned from posting, and each interaction is converted into a feature vector.
USLegis:11footnotemark: 1 USLegis is a Senate co-sponsorship network. It records instances where members of the United States Senate co-sponsor the same bill in a given Congress. The dataset has 225 nodes and 60,396 edges.
Wikipedia:22footnotemark: 2 This dataset contains 157,474 temporal edges with 1,000 most edited pages and 8227 active users as nodes. Labels indicate whether users are prohibited from editing the page, and the interactions are transformed into text features. Both Reddit and Wikipedia are bipartite graphs.
ComplexDynamics: [^3] https://www.dgl.ai/WSDM2022-Challenge/ This dataset is a dynamic interaction graph with entities as nodes and different types of events as edges, a non-bipartite graph as the training dataset of the WSDM 2022 Challenge. To generate this dataset, we choose 8295 different nodes, and the number of interactions per node is less than 753, resulting in 150035 interactions.
To generate the Complex Dynamics dataset, we started with the snapshot of the WSDM 2022 Challenge training data on 2014-10-19 04:00:00. We first cleaned the data by filtering out all missing and unavailable data. Then, in the remaining data, we kept only nodes with less than 753 interactions where the timestamp (the timestamp in Unix epoch) was present. The data processed in this way would cause confusion in the node sequence number and time sequence, so we had to renumber the nodes (with anonymized categorical features) and sort each interaction event (with anonymized categorical features) under the time constraint, resulting in 150035 interactions generated by 8295 nodes. The resulting dataset has the following characteristics: 1) the number of high-frequency interactions between nodes is relatively small; 2) the rate of continuously repeated interactions between the same nodes or node pairs is low; 3) there are symbiotic events under the same time stamp.
TABLE I: Results of the transductive link prediction task on MOOC, UNvote and USLegis, with AP, AUC and ACC metrics, highlighting the best-performing results for each dataset in bold. All results are shown as percentages, obtained by multiplying the values by 100 for standardization.
Dataset MOOC UNVote USLegis
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 89.37±1.6 91.17±1.3 84.13±1.5 66.13±1.2 70.31±1.9 63.97±0.9 73.91±0.7 81.37±0.7 73.71±0.6
GraphSAGE-t 84.76±1.5 87.31±1.4 80.37±1.3 65.76±0.9 70.72±1.1 63.07±0.7 71.98±0.8 78.32±0.9 71.20±0.9
JODIE 80.47±2.0 83.61±1.6 76.59±1.9 62.99±1.1 67.65±1.2 62.43±1.0 73.31±0.4 81.57±0.2 73.39±0.2
TGAT 85.09±0.4 86.20±0.4 77.40±0.4 51.22±0.6 51.49±0.9 51.07±0.7 65.75±6.0 72.55±7.8 65.16±5.1
DyRep 77.73±3.9 81.41±3.6 74.50±3.4 62.38±0.6 66.81±0.7 61.80±0.5 69.16±3.0 76.68±3.3 69.53±2.6
TGN 88.89±1.5 90.89±1.2 83.29±1.5 65.62±1.2 70.11±1.4 64.16±1.0 75.13±1.3 82.39±1.0 74.31±0.8
CAWN 69.15±0.3 70.20±0.3 63.44±0.2 51.71±0.1 52.75±0.1 50.76±0.2 69.94±0.4 75.83±0.5 66.92±0.5
EdgeBank 52.91±0.0 54.84±0.0 50.08±0.0 54.70±0.0 58.34±0.0 58.34±0.0 55.45±0.0 59.03±0.0 56.48±0.0
TCL 82.60±0.8 83.15±0.5 74.31±0.4 50.81±0.7 50.91±1.0 50.58±0.6 68.54±1.2 75.49±0.7 67.06±0.8
GraphMixer 82.00±0.2 83.31±0.2 74.56±0.1 52.28±0.5 53.31±0.5 51.92±0.3 69.28±0.9 76.09±0.6 67.40±0.7
DyGFormer 86.91±0.1 86.53±0.1 77.32±0.5 54.09±0.2 54.94±0.4 53.37±0.3 68.31±1.0 75.52±0.8 67.35±1.0
MPFA 92.98±0.3 94.46±0.2 88.09±0.3 67.39±2.1 71.60±2.3 65.51±1.9 76.83±0.3 84.26±0.1 76.46±0.2
TABLE II: Transductive link prediction task results on LastFM, SocialEvo and ComplexDynamics using AP, AUC and ACC metrics, with the best results for each dataset highlighted in bold. All results are presented as percentages, obtained by multiplying the values by 100 for standardization.
Dataset LastFM SocialEvo ComplexDynamics
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 70.37±1.7 71.29±1.3 67.46±1.3 90.33±0.2 91.36±0.9 87.41±1.5 88.12±0.1 87.04±0.1 78.54±0.2
GraphSAGE-t 73.21±1.8 74.32±1.3 67.76±0.9 90.21±1.5 92.59±0.9 86.29±1.3 86.95±0.2 85.67±0.3 77.34±0.4
JODIE 70.21±1.9 69.88±1.4 64.83±0.9 88.55±1.8 90.96±1.3 83.77±1.5 82.93±0.8 82.67±0.6 74.59±0.6
TGAT 72.43±0.3 70.69±0.2 64.52±0.2 92.71±0.7 94.32±0.5 89.32±0.3 84.57±0.2 83.01±0.1 74.49±0.2
DyRep 70.55±2.1 70.17±1.9 65.00±1.3 85.11±0.5 89.05±0.3 82.13±0.2 79.17±0.7 78.77±0.4 71.66±0.3
TGN 75.29±3.7 76.41±3.2 69.91±2.6 93.70±0.3 95.68±0.2 90.32±0.4 87.84±0.3 86.82±0.2 78.35±0.3
CAWN 73.36±0.2 69.43±0.1 63.48±0.2 78.46±0.3 80.43±0.4 74.89±0.6 90.25±0.1 88.35±0.2 79.39±0.2
EdgeBank 77.25±0.0 83.51±0.0 82.06±0.0 52.04±0.0 53.90±0.0 53.90±0.0 53.90±0.0 70.82±0.0 50.00±0.0
TCL 69.39±2.6 64.67±2.1 59.47±1.7 93.07±0.2 94.38±0.2 90.22±0.1 89.18±0.1 87.30±0.1 78.32±0.1
GraphMixer 75.68±0.2 73.42±0.3 67.21±0.2 92.56±0.4 94.47±0.3 89.35±0.2 87.04±0.2 85.79±0.2 77.50±0.1
DyGFormer 83.81±0.2 79.97±0.2 71.63±0.3 94.21±0.1 95.56±0.2 91.83±0.2 89.28±0.1 86.84±0.1 77.55±0.1
MPFA 84.23±0.8 84.78±0.9 76.99±0.6 95.75±0.3 93.72±0.2 90.55±0.2 89.93±0.3 87.84±0.3 79.91±0.4
TABLE III: Results of the inductive link prediction task on MOOC, UNvote and USLegis, with AP, AUC and ACC metrics, highlighting the best-performing results for each dataset in bold. All results are shown as percentages, obtained by multiplying the values by 100 for standardization.
Dataset MOOC UNVote USLegis
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 87.36±2.5 88.76±2.4 81.32±2.7 57.71±3.4 59.21±4.6 55.73±3.6 59.16±1.7 60.93±1.3 57.78±0.7
GraphSAGE-t 84.23±1.5 86.72±1.2 78.32±1.1 57.65±1.3 58.97±1.4 55.37±1.9 60.31±0.7 59.37±0.4 57.31±0.6
JODIE 80.54±0.9 83.44±0.9 76.43±0.9 55.74±1.7 56.65±2.7 53.58±2.12 52.16±0.5 55.07±0.8 53.15±0.5
TGAT 84.53±0.5 85.68±0.4 77.02±0.5 52.21±1.0 51.96±1.1 51.61±0.76 49.71±2.3 47.33±3.5 48.68±2.9
DyRep 78.52±2.8 82.11±2.3 74.99±2.2 54.38±1.6 54.77±2.3 52.37±1.81 56.26±2.0 58.87±2.1 55.56±1.6
TGN 88.79±1.4 90.82±1.3 83.25±1.5 57.14±2.6 58.90±4.0 55.28±3.00 59.52±1.3 62.14±1.7 58.26±1.0
CAWN 68.27±0.6 68.66±0.7 61.91±0.6 49.74±0.7 48.32±0.6 47.67±0.33 53.11±0.4 51.01±0.5 50.11±0.8
TCL 80.95±0.8 81.57±0.5 72.98±0.4 50.20±1.0 49.90±1.5 49.32±1.16 49.43±1.0 46.84±1.1 48.19±1.0
GraphMixer 80.56±0.2 81.99±0.1 73.22±0.2 51.01±0.8 50.97±0.9 50.17±0.7 49.31±1.5 46.02±1.3 47.23±1.3
DyGFormer 85.66±0.2 85.20±0.2 76.05±0.2 53.87±0.2 54.31±0.2 53.16±0.24 48.08±0.6 45.34±0.9 46.84±1.0
MPFA 92.57±0.5 94.19±0.4 87.67±0.6 63.87±4.0 65.76±4.1 61.24±3.3 60.76±1.3 62.67±1.2 58.61±0.8
TABLE IV: Inductive link prediction task results on LastFM, SocialEvo and ComplexDynamics using AP, AUC and ACC metrics, with the best results for each dataset highlighted in bold. All results are presented as percentages, obtained by multiplying the values by 100 for standardization.
Dataset LastFM SocialEvo ComplexDynamics
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 79.73±2.1 80.32±1.7 73.18±1.1 82.33±0.2 83.21±3.1 76.36±3.1 80.15±0.3 78.49±0.3 70.75±0.4
GraphSAGE-t 81.32±1.3 81.29±0.7 75.63±0.5 79.66±5.7 83.57±4.5 77.23±4.0 80.26±0.7 78.37±0.8 70.73±0.6
JODIE 82.36±1.1 81.63±1.0 75.34±0.8 91.22±1.9 93.10±1.4 87.68±1.8 73.90±1.0 71.63±1.0 65.44±0.7
TGAT 77.74±0.4 76.27±0.2 69.90±0.8 91.25±0.4 93.10±0.8 87.54±1.0 83.00±0.3 80.87±0.3 72.76±0.4
DyRep 81.65±1.8 81.01±1.9 74.75±1.4 82.37±0.6 84.87±0.3 78.56±0.2 67.80±1.1 66.77±1.0 61.73±0.7
TGN 80.58±3.8 81.51±3.0 75.26±2.6 90.70±0.7 93.37±0.4 89.32±0.5 79.40±0.6 77.53±0.6 69.57±0.4
CAWN 77.36±0.3 73.16±0.4 67.43±0.3 77.17±0.3 79.46±0.4 74.67±1.0 87.37±0.1 84.28±0.2 75.10±0.4
TCL 72.07±3.2 68.76±2.9 63.21±2.7 91.97±0.3 93.13±0.2 89.53±0.2 87.86±0.1 85.31±0.1 76.53±0.3
GraphMixer 82.14±0.4 80.23±0.3 73.99±0.2 91.23±0.4 93.28±0.3 88.13±0.2 80.95±0.4 78.24±0.4 70.84±0.3
DyGFormer 86.96±0.2 83.82±0.3 76.17±0.3 92.26±0.1 94.32±0.2 90.67±0.2 88.25±0.1 85.13±0.1 76.01±0.3
MPFA 88.05±0.7 88.55±0.5 81.60±0.7 94.65±0.2 92.32±0.2 90.36±0.3 84.35±0.5 83.64±0.2 73.48±0.3
TABLE V: Dynamic node classification with AUC. All results have been converted to percentages by multiplying by 100.
Wikipedia Reddit
GAT 82.34±0.8 64.52±0.5
GraphSAGE 82.42±0.7 61.24±0.6
CTDNE 75.89±0.5 59.43±0.6
JODIE 84.84±1.2 61.83±2.7
TGAT 83.69±0.7 65.56±0.7
DyRep 84.59±2.2 62.91±2.4
TGN 87.81±0.3 67.06±0.9
CAWN 84.88±1.3 66.34±1.8
TCL 77.83±2.1 68.87±2.2
GraphMixer 86.80±0.8 64.22±3.3
DyGFormer 87.44±1.1 68.00±1.7
MPFA 90.67±0.2 72.33±0.5
V-B Baselines
We select nine temporal models (JODIE [24], DyRep [21], TGAT [26], TGN [25], CAWN[29], EdgeBank[30], TCL[27], GraphMixer[39], and DyGFormer[28]) as benchmarks. Additionally, we enhance two superior static graph models, GAT and GraphSAGE, to temporal models (GAT-t and GraphSAGE-t) and use them as new baselines. Refer to the Related Work subsection for baseline details. To ensure fair comparisons, all baselines share consistent settings across eight datasets: batch size (200), number of temporal neighbors (10), embedding dimension (172), and dropout rate chosen from (0.0, 0.1, 0.2, 0.3, 0.4). The learning rate is set to 0.0001. The remaining baseline settings align with those specified in their respective papers. The extension of GAT and GraphSAGE to temporal models involves incorporating temporal information.
V-C Experimental setup and evaluation metrics
For fair comparisons, our model and baselines share identical hyperparameters whenever possible. Across all datasets and experiments, we maintain a batch size of 200, a learning rate of 0.0001, and an embedding dimension of 172. In the link prediction task, we run experiments on six temporal datasets. During training, we compute event probabilities using an equal number of negative and positive samples. Reference metrics include average precision (AP), accuracy (ACC), and area under the ROC curve (AUC). For the node classification task on the benchmark datasets (Wikipedia and Reddit), AUC is employed due to label imbalances. Temporal neighbors, selected from the 10 nearest neighbors in chronological order, are used in all experiments. The final results are the average of 10 model runs. For all tasks, we split each dataset chronologically into training (75%), validation (15%), and test (15%) sets.
V-D Experiments
V-D1 Dynamic link prediction
We conduct experiments on six temporal datasets, i.e. USLegis, ComplexDynamics, SocialEvo, Mooc, UNVote, LastFM for dynamic link prediction. The task is to determine, given nodes $v_{i}$ and $v_{j}$ at time $t$, whether they will interact with each other? Within this task, we design two additional subtasks, the inductive and the transductive subtasks. The transductive task indicates that the nodes appearing in the test set were seen by the model during training, while they were not seen in the inductive task.
From the results of dynamic link prediction on the six datasets (see Tables I, II, III, and IV), our proposed algorithm demonstrates a distinct advantage in this task. Across four datasets—MOOC, UNvote, USLegis, and LastFM—MPFA consistently outperforms other algorithms in both transductive and inductive subtasks based on evaluation criteria such as AP, AUC, and ACC. For instance, in the transductive task on the MOOC dataset (see Table I), MPFA achieves an AP score of 92.98, surpassing the second-best algorithm (GAT-t) by 3.61 percentage points and outperforming classical algorithms TGAT and TGN by 7.9 and 4.09 percentage points, respectively. MPFA also exhibits advantages in SocialEvo and ComplexDynamics, particularly in the AP scoring criterion (refer to Table II). However, on SocialEvo and ComplexDynamics, MPFA slightly lags in ACC and AUC. This discrepancy may be attributed to a frequency dominance feature in these datasets, favoring algorithms like DyGFormer and TCL. Examining the two subtasks, both MPFA and the baselines demonstrate a stronger advantage in the transductive task across most datasets. This is primarily due to the fact that, in the transductive task, the model predicts nodes seen during training, whereas the inductive task involves predicting unseen nodes. An interesting anomaly is observed in the LastFM dataset, where the inductive task outperforms the transductive task. Taking MPFA as an example, the AP, AUC, and ACC for the inductive task are 88.05, 88.55, and 81.60, respectively, while the corresponding values for the transductive task are only 84.23, 84.78, and 76.99 (refer to Table II and IV). In summary, our proposed continuous-time dynamic graph modeling algorithm proves to be effective and efficient in predicting dynamic links.
(a) Inductive results
(b) Transductive results
(c) Inductive results
(d) Transductive results
Figure 4: Ablation experiment results. Results of MOOC ablation experiments are denoted by (a) and (b), and for the USLegis dataset by (c) and (d). W/O RP: without Raw Perspective; W/O EP: without Evolving Perspective; W/O RED: Raw and Evolving perspectives without Dynamic update; W/O ED: Evolving perspective without Dynamic update.
(a) MOOC
(b) USLegis
Figure 5: Dynamic link prediction results for six models with varying numbers of history neighbors on MOOC and USLegis reveal that superior predictions with fewer history neighbors suggest the model requires only a limited number of interactions for effective long-term dependency properties.
(a) MOOC batch
(b) MOOC embed
(c) MOOC neighbors
Figure 6: Hyperparameter experiment results on the MOOC dataset.
(a) USLegis batch
(b) USLegis embed
(c) USLegis neighbors
Figure 7: Hyperparameter experiment results on the USLegis dataset.
(a) Inductive results
(b) Transductive results
(c) Inductive results
(d) Transductive results
Figure 8: Visualization of attention coefficients for transductive and inductive tasks on the MOOC dataset is presented in (a) and (b), while (c) and (d) represent the results for the USLegis dataset.
(a) MOOC
(b) USLegis
Figure 9: Model runtime on the MOOC dataset is denoted as (a), and on the USLegis dataset as (b).
V-D2 Dynamic node classification
We perform the dynamic classification experiment on two datasets with dynamic labels (Wikipedia and Reddit). The purpose of this experiment is to determine whether a node is banned after going through an event. For example, on Wikipedia, when a user posts, the platform determines whether to ban them based on their posts. After obtaining the temporal network embeddings, we train an MLP decoder to predict node label changes. Due to the label imbalance in the dataset, we only use AUC as the evaluation metric in this experiment. For dynamic node classification, Table V compares the results of MPFA and baselines, where our method achieves state-of-the-art results on both benchmark datasets. We outperform the second best method by 2.9 and 4.3 on the Wikipedia and Reddit datasets, respectively. Thus, our proposed model is also valid for the dynamic classification task.
V-D3 Ablation studies
To evaluate the effectiveness of different components of our proposed model, we conduct the ablation studies in this part. We experiment with different combinations of modules from MPFA for dynamic link prediction on the MOOC and USLegis datasets, aiming to observe their specific properties (see Figure 4 for details). From the results of the overall ablation experiments on both datasets, MPFA with two perspectives achieves the highest performance in both transductive and inductive tasks, as evidenced by superior values in AP, AUC, and ACC. This result substantiates the effectiveness of MPFA. Conversely, the evolving perspective without dynamic updating emerges as the least effective, as indicated by the lowest values across all three evaluation criteria. This underscores the significance of the dynamic updating module designed to capture long-term dependencies. Regarding the evolving and raw perspectives, the model exclusively featuring the evolving perspective demonstrates the second-best performance on MOOC, while on USLegis, the model with only the raw perspective achieves the second-best performance. This discrepancy highlights the varying roles of these perspectives in different scenarios. Furthermore, the model incorporating both evolving and raw perspectives but lacking dynamic updating performs suboptimally, underscoring the crucial role of long-term dependency in effective model learning.
V-D4 Long dependency propertity
To verify that our model can maintain the long-term dependency property with a smaller number of historical neighbors, we conduct experiments on link prediction performance under different numbers of temporal neighbors for MPFA and five baselines on the MOOC and USLegis datasets. As shown in Figure 5, we observe that MPFA maintains a stable state when the number of temporal neighbors varies from 1 to 30, both on the MOOC and USLegis datasets. In comparison to other models, especially on the MOOC dataset, all five baselines exhibit the characteristic of performance improvement with an increasing number of neighbors. Among them, CAWN shows the most significant performance improvement. Although these baselines occasionally display jumps in predicted values for the transductive and inductive subtasks of USLegis, they still require an increase in the number of neighbors to achieve performance improvement. In summary, our proposed dynamic update module is effective. Superior predictions with fewer history neighbors suggest that our model requires only a limited number of interactions for effective long-term dependency properties.
V-D5 Hyperparametric analysis
In this section, we examine the performance of MPFA under three main hyperparameters on two datasets, MOOC and USLegis. As illustrated in Figure 6, MPFA’s performance remains relatively stable with variations in batch size, embedding dimension, and the number of temporal neighbors. In Figure 6(a), the model performance exhibits a slight decrease as the batch size increases. In Figure 6(b), there is a slight increasing trend in model performance with higher embedding dimensions. Figure 6(c) shows the model remains stable up to 30, with a slight decrease thereafter. In USLegis, MPFA experiences two significant fluctuation points with different batch sizes: 200 points in the transductive task and 300 points in the inductive task (Figure 7(a)). Across varying embedding dimensions (Figure 7(b)), similar to the MOOC dataset, MPFA’s performance tends to increase slightly. In experiments with different numbers of temporal neighbors (Figure 7(c)), performance in the transductive task remains stable, with some fluctuations observed in the inductive task.
V-D6 Attention score visualization
In this section, we visualize attention coefficients for both perspectives on the MOOC and USLegis datasets (see Figure 8). This enables observation of the focus areas for the two perspectives in transductive and inductive tasks. Generally, the evolving perspective exhibits higher attention coefficients, typically reaching values of 0.3 or higher, while the raw perspective tends to have smaller coefficients, usually below 0.4. These coefficient values align with the model’s ability to gain deeper insights into historical events post-learning, emphasizing extraction of a limited amount of valid information. On both the MOOC and USLegis datasets, the evolving perspective in the transductive task directs more attention to neighbors closer in the time series. Specifically, on the MOOC dataset, the evolving perspective exhibits exceptional attention to the nearest temporal neighbors. Conversely, in the inductive task, the focus is relatively more diffuse. For the raw perspective on both datasets, temporal attention on the MOOC dataset is notably larger and broader in the transductive task. On the USLegis dataset, attention remains essentially the same for both transductive and inductive tasks.
V-D7 Time efficiency
To assess the runtime efficiency of MPFA and current leading baselines, we select five algorithms (TCL, TGAT, GraphMixer, CAWN, DyGFormer) to evaluate the time consumption for achieving optimal training results on the MOOC and USLegis datasets, as depicted in Figure 9. Although TGAT requires the least time to optimize its model parameters on MOOC, it does not perform well in terms of link prediction results. In contrast, MPFA, despite consuming more time than TGAT, achieves optimal test performance on MOOC. Similarly, TCL also has the shortest training time but fails to yield better test results. We attribute this to both models potentially inadequately fitting the dataset, causing premature training termination. Overall, MPFA demonstrates an advantageous balance between runtime efficiency and optimal performance.
V-E Discussion
Graph neural networks have proven highly effective for processing graph-structured data, with numerous algorithms proposed for various scenarios of static graphs. However, in real-world applications, network structures are not static but constantly evolve over time. Consequently, conventional static graph expression algorithms struggle to adapt to these dynamically changing scenarios. Several models have been introduced to address dynamic graph problems, categorized into discrete and continuous models. Discrete models, which handle network evolution through snapshots extracted at specific time intervals, suffer from information loss due to their event stream representation. Continuous dynamic graph models, while capturing the moment of evolution, often neglect the raw interaction state, diminishing the impact of historical events on predicting future events. Moreover, many current continuous-time dynamical graph models excessively consume time resources to obtain long-term dependencies. To overcome these challenges, this paper proposes distinct treatment schemes:
Handling Changing Topology: For static graphs unable to adapt to changing topology, we transform the dynamic graph network into an event stream, creating a fine-grained model for comprehensive consideration of network evolution details. Simultaneously, we assess whether classical static models maintain an advantage when extended to temporal graphs by adapting GAT and GraphSAGE. Despite performance differences compared to MPFA, experimental results show they outperform many dynamic models.
Incorporating Multiple Perspectives: To address the limitation of existing continuous-time dynamic models focusing solely on the evolving perspective, we propose incorporating a raw perspective to observe the network’s evolution from multiple directions. Ablation experiment results highlight distinct roles for each perspective in various scenarios. By coupling the evolving and raw perspectives, we enhance the model’s applicability to different contexts.
Dynamic Updating Model for Long-Term Dependencies: Existing models typically achieve long-term dependencies by increasing the number of historical neighbors. In contrast, we adopt a dynamic updating model, allowing the use of a small number of historical neighbors to acquire long-term dependencies.
Despite achieving good results, our model faces certain challenges. For instance, when handling networks with a small number of nodes (a few hundred nodes) and millions of edges, its performance tends to be slightly lower than some contemporary models. Additionally, increasing the batch size to a large value results in a decrease in model performance. These issues need to be addressed in future work.
VI Conclusion
In this paper, we address dynamically changing graph networks through event streams, proposing an effective continuous-time dynamic graph model. Our model pioneers the exploration of dynamic network evolution from both the raw and evolving perspectives. The evolving perspective captures real-time network dynamics, revealing fine-grained topology transformations and long-term dependencies with fewer historical interactions. Ablation experiments and long-term dependency characterization confirm the effectiveness of the evolving perspective. The often-overlooked raw perspective, representing the nature of interaction events, shows its advantages in diverse datasets or domains through ablation experiments. Relying on only one perspective may weaken the model’s generalization ability, emphasizing the importance of simultaneous learning from both perspectives. Validation through dynamic link prediction, dynamic classification, and numerous experiments underscores the effectiveness of our proposed model. This approach introduces a novel idea for future research and holds promise for application across multiple scenarios with dynamic changes.
Acknowledgments
The authors would like to thank all the experts who helped with this project and suggested revisions to the paper, as well as those who provided open source code and public datasets. At the same time, we would like to thank colleagues and experts for taking the time out of their busy schedules to provide valuable opinions.